Shift report 20011107 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 07 2001 - 07:50:10 EST

  • Next message: Flemming Videbaek: "Re: Shift report 20011106 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Nov  7 07:50:09 2001, Shift: 08:00-16:00
    Supervisor: BHS, with: 
    Log book#10 pages 108 - 110
    
    Shift Summary:
    ------------------------------------------------------------
    1: A store had just been made when the shift began. This lasted until 03:50, when it was dumped by a BRAHMS/PHENIX vote (as always). After this they had a few QLIs and then started beam studies at 06:00, which is still going on.
    
    2: Data runs:
    Run      Trig 6     Settings
    5701     116k      MRS@90, D5@4A
                             FFS@20, BFS@6, DCs OFF, D1-D4@1/10A
    5702     111k      Same
    5703       12k      Same, except D5@2A (350 amps)
    --------------------------------------
                229k trigger 6
    
    After this I turned D5 off for a 0-field run becaus MCR had promised another store, but no such luck.
    5704    Pedestal run
    5705    TPC pedestal run
    5706    DVM pedestal run 
    
    3: No detector problems. A printer problem though - the SuperMon plots tend to jam up the queue for qms_ch and have to be killed. Is this something that can be fixed? Also, I was unable to find Tpcped - could someone who knows where it is put a link ( ! ;-) to it on the operator desktop? Consequently the results of my TPC pedestal run are not installed yet.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 07 2001 - 07:50:38 EST